OpenStudy (anonymous):

The circumference of a circle is 47.1 feet. What is the diameter of the circle?

OpenStudy (ipwnbunnies):

For any circle, \[Circumference = \pi*diameter\]

OpenStudy (ipwnbunnies):

Therefore, \[diameter = \frac{circumference}{\pi}\]